On a spacecraft, two engines are turned on for 482 s at a moment when the velocity of the craft has x and y components of v0x = 3460 m/s and v0y = 2670 m/s. While the engines are firing the craft undergoes a displacement that has components of x = 2.11×106 m and y = 3.07×106 m. Find the x and y components of the craft’s acceleration.
1
Expert's answer
2020-02-05T12:13:40-0500
"x=v_{ox}t+0.5a_xt^2"

"y=v_{oy}t+0.5a_yt^2"

"2.11\cdot 10^6=(3460)(482)+0.5a_x(482)^2"

"3.07\cdot 10^6=(2670)(482)+0.5a_y(482)^2"

"a_x=3.81\frac{m}{s^2}"

"a_y=15.3\frac{m}{s^2}"